OpenStudy (honeymonster):

how do i make x the subject of the formula with this equation. (x-1)(y+2)=3

hartnn (hartnn):

to isolate x,see what is multiplied with it so when u divide that on both sides,u can isolate x, so what should be divided?

OpenStudy (honeymonster):

i've multiplied the brackets out and i've come up with xy+2x-y-2=3..so it would be divided by 2 or y?

hartnn (hartnn):

ok,since u have already multiplied the brackets,now u don't need to divide anything. Now u take all the terms which do not contain x on other side. like here u have to take -y and -2 on other side,u know how u can do that??

OpenStudy (honeymonster):

i've tried to rearrange it.. xy+2x =3+y+2..i got this

hartnn (hartnn):

correct :) now take x common from left side from xy+2x then what u get? remember : ab+ac=a(b+c)

OpenStudy (amistre64):

oy thats alot of extra work ;)

hartnn (hartnn):

yes,but as u can see, honeymonster has already multiplied the brackets.....

OpenStudy (amistre64):

have him put them back together ... oh the humanity!

hartnn (hartnn):

its almost done,taking common and then dividing it......

OpenStudy (honeymonster):

would you divide it by y+2?

OpenStudy (amistre64):

to begin with, that would be a swell idea

OpenStudy (amistre64):

\[(x-1)(y+2)=3~:~divide~by~(y+2)\] \[\frac{(x-1)\cancel{(y+2)}}{\cancel{(y+2)}}=\frac{3}{(y+2)}\] \[x-1=\frac{3}{(y+2)}\]

OpenStudy (honeymonster):

ok so then how do you get rid of the -1?

OpenStudy (amistre64):

if you owe someone 1 dollar, how do you get to the point where you owe them nothing?

OpenStudy (honeymonster):

give it back..so you would add it to the otherside?

OpenStudy (amistre64):

yes; add 1 to both sides

OpenStudy (amistre64):

\[x-1+1=\frac{3}{(y+2)}+1\] \[x-0=\frac{3}{(y+2)}+1\] \[x=\frac{3}{(y+2)}+1\]

OpenStudy (honeymonster):

ok thats great! thank you for the help :)

OpenStudy (amistre64):

youre welcome, it was a communal effort :)